OpenStudy (anonymous):

Five of Ethan's family members went on a 6 1/4 road trip. If they all shared the drive time equally, how many hours did each family member drive? 4/5, 1 1/4, or 1 3/20

OpenStudy (phi):

they want you to divide 5 into \( 6 \frac{1}{4}\) I would change \( 6 \frac{1}{4}\) into an improper fraction (do you know how?) and then multiply by \(\frac{1}{5} \)

OpenStudy (anonymous):